# Compositional Difference-in-Differences for categorical outcomes

### v1

@@ 1. Canonical $2\times 2$ DiD Case
In this section, we focus on the canonical, two-group, two-time period case in the DiD framework, and we clearly expose the assumptions and their implications. We will start with the analytical framework, later discuss the main identifying assumption, and all the economic and geometric implications.
1.1. Analytical framework
In this section, we introduce the basic framework of our method. A categorical outcome $Y$ takes values in a finite set of $p$ mutually exclusive categories:
@@ 1.1. Analytical framework
We focus primarily on the absolute quantities for each category, which can appear in several empirical contexts. They can arise when the outcome is observed at the individual level. For example, in evaluating a job training program, individuals may be classified as unemployed, part-time employed, or full-time employed. The categorical quantity is obtained by aggregating individuals into these outcome categories. Sometimes, the data consist directly of quantities across categories rather than individual-level observations. For example, the number of votes each political party receives in an election. Unlike individual-level data, the identities or attributes of the units are not observed. In all these cases, the proposed method remains applicable, as it relies solely on knowledge of the categorical quantities.
We consider a setting with two groups, $g\in\{0,1\}$, where $g=1$ is the treated group and $g=0$ is the control group. These groups are observed in two time periods, $t\in\{0,1\}$, with $t=0$ being the pre-treatment period and $t=1$ the post-treatment period. Our object of interest is a categorical outcome. For any category $c_{k}$, we define $q^{N}_{g,t}(c_{k})$: the potential untreated quantity (the quantity in $c_{k}$ if the group $g$ were untreated at time $t$). We also have $q^{I}_{g,t}(c_{k})$: the potential treated quantity (the quantity in $c_{k}$ if the group $g$ were treated at time $t$). For the untreated potential outcomes, we aggregate across categories to define the total quantity for a group-period:
$S^{N}_{g,t}=\sum_{k=1}^{p}q^{N}_{g,t}(c_{k}).$
@@
All the same definitions apply to the treated potential quantities, $q^{I}_{g,t}(c_{k})$, and their associated totals and shares. From the data, we can identify the vectors $q^{N}_{0,0}$ for the control group and $q^{N}_{1,0}$ for the treated group in the pre-treatment period. In the post-treatment period, we can identify from data the untreated vector of quantities for the control group, $q^{N}_{0,1}$, and the treated counterpart for the treated group, $q^{I}_{1,1}$. The counterfactual of interest is the vector of quantities that we would have observed for the treated group in the absence of treatment $q^{N}_{1,1}$. Identifying this missing object is the core challenge of our analysis. Once this is identified, we can consider the treatment effect parameters.
We introduce two treatment effect parameters to capture the impact of the intervention. The first is the growth treatment effect on the treated (GTT), which quantifies the causal effect on the absolute size of a category. For each category $c_{k}$ (where $k=1,\ldots,q$), the GTT is defined as the proportional change in its quantity:
@@
The second parameter is the compositional treatment effect on the treated (CTT), which captures how treatment redistributes weight across categories. To measure this, we apply the compositional difference operator commonly used in compositional data analysis to compare probability mass functions. This operator quantifies how each category expands or contracts relative to all others. For intuition, fix a category $c_{k}$ and consider its proportional change:
@@
Dividing by the total change gives a vector showing each category’s relative gain or loss. We obtain the compositional difference operation used in compositional data analysis for the control group.
@@
1.2. parallel growths and Identification
This section develops the identification strategy and clarifies its implications for the probability mass function. We start with the following assumption.
@@
This assumption ensures that all observed and counterfactual outcomes in the absence of treatment are defined over a common set of categories. It rules out situations where the support changes across groups or over time, which would make comparisons and counterfactual reasoning impossible. In practice, it implies that the set of possible outcomes is fixed. Such support conditions are standard in the literature. To introduce our main identifying assumption, we provide intuition for the DiD identification logic. The core idea is that, in the absence of treatment, the treated group would have followed the same trajectory as the control group. Figure 2 illustrates this DiD identification logic: the control group, unexposed to the treatment, serves as a counterfactual benchmark, mirroring what would have happened to the treated group had it not received the intervention.
@@
As we said before, in many empirical settings, the sizes of the groups defined in each category can vary dramatically in scale. Therefore, a standard parallel trends assumption on the raw quantities is problematic in this context, as it would impose the same absolute change on both large and small groups, which is often unrealistic. To address this issue of scale and to formulate a more plausible identifying assumption, we instead consider parallel trends on the log transformation of the quantities. This approach focuses on proportional, rather than absolute, changes. For a given vector, define the functions $(\log\cdot)$ and $(\exp\cdot)$ which apply respectively the logarithm and the exponential functions component-wise to the vectors.
@@
Assumption 2 is the core identifying condition for our model. It states that, in the absence of treatment, the proportional growth (or decline) of each category in the treated group would have been the same as that of the control group. This reflects the intuitive idea that underlying trends affect all categories proportionally to their size. This may appear as a more realistic and flexible assumption than one based on raw quantities in many economic contexts, especially when the sizes of categories differ significantly. It captures the idea that trends often affect all units proportionally to their size. It is also possible to use discrete covariates to make the assumption more believable. We discuss this possibility in appendix B. From this assumption, the counterfactual quantity vector for the treated group can be easily recovered, and the corresponding counterfactual probability mass can be obtained by normalization, as we see in Theorem 1. Taken together, assumptions 1 and 2 impose sufficient structure to point-identify the counterfactual distributions. The following theorem 1 formalizes this result
@@
This identification result is constructive and provides a straightforward way to implement the method. In the next section, we want to discuss further justifications for parallel growths using its economic and geometric implications.
1.3. Implications of parallel growths
Assumption 2 restricts the evolution of growth rates of quantities but also of shares. In this section, we give further economic and geometric justifications for parallel growths.
Interpretation using the random utility model: To formalize that, let’s define the multinomial logistic transform. For $\pi=[\pi(c_{1}),\dots,\pi(c_{p})]\in\mathcal{S}^{p-1}$. The log-odds (or multinomial logit) transformation $\ell:\mathcal{S}^{p-1}\to\mathbb{R}^{p-1}$ maps probabilities to real-valued indices:
@@ 1.3. Implications of parallel growths
The parallel growths assumption would imply a parallel trends assumption on the log-transformation of the probability distribution, as stated in this proposition.
@@
Assumption 1, and assumption 2 implies the parallel trends of log-odds, which is: $\ell(\pi_{1,1})-\ell(\pi_{1,0})=\ell(\pi_{0,1})-\ell(\pi_{0,0}).$
This implication is important because it helps to better formalize the economic intuition behind our main assumption. In fact, it implies that, in the absence of treatment, the change in log-odds of each category (relative to a base category) would be the same across the treatment and control groups. To better explain the value of this implication, we analyze it through the lens of the random utility model (McFadden, 1972, 1974). In this foundational economic framework, individuals choose the option that provides the highest unobserved utility. The log-odds of choosing one category over another directly measure the difference in their underlying utilities (see appendix D), providing a direct measure of the underlying preference structure driving people’s choices. For example, in a voting context, the log-odds of choosing Democrat over Other (third parties) quantifies the relative preference for the Democratic party. Our parallel growths assumption, when viewed through this lens, implies that the relative preferences between any two categories would have evolved in parallel for the treated and control groups in the absence of the treatment. To illustrate: suppose the control group shows an increase in the log-odds of voting Democrat versus Other, indicating a growing preference for Democrats over third parties. Our assumption states that the treated group would have experienced this same evolution in relative preferences had they not been subjected to the treatment. This same logic applies to all category pairs, such as Republicans versus Other. This interpretation is important because it grounds our technical parallel growths condition to a clear behavioral story. It is no longer just an abstract statistical constraint, but an economic hypothesis about the stability of relative preference trends across groups through the random utility model. From that, we can also identify the counterfactual distribution with an alternative formula:
@@
Equation (8) represents a compositional DiD adjustment in the transformed space: the treatment-free evolution of the distribution is added to the initial treated distribution, and the result is mapped back to the simplex through the inverse transformation $\ell^{-1}(\cdot)$. This formulation not only preserves the probabilistic structure of the outcome (ensuring that $\pi_{1,1}^{N}$ remains a valid probability vector) but also provides an intuitive and computationally convenient way to estimate the counterfactual distribution when dealing with categorical or compositional outcomes.
Interpretation using the compositional difference interpretation: To further support the relevance of our assumption in this context, we provide an alternative interpretation that does not depend on the random utility framework or its underlying assumptions. Suppose our goal is to understand how the treatment redistributes probability mass across categories, that is, how each category gains or loses relative shares compared to the others as a consequence of treatment. An intuitive way of building the counterfactual would be to assume that the relative growth/decline in shares of each category in the control group reflects what would have happened in the treated group without treatment. The compositional difference operator $(\ominus)$ defined in equation 1 provides the precise mathematical language to formalize this intuition. We now show that our parallel growths assumption can be interpreted using that intuition. Specifically, assumption 2 implies that, in the absence of treatment, the relative growth or decline of category shares would have been similar across the treated and control groups. This interpretation is formally established in the following proposition.
@@
In figure 3, each point in the triangle represents a distribution over three categories. As time evolves, distributions trace curves within the simplex. Although these paths may look non-linear in Euclidean space, in the simplex geometry, they are parallel, maintaining a consistent direction without intersecting. Here, we say that the implied evolution of the distributions is consistent with the nature of their space. Because of that, the compositional treatment effect on the treated can be written as:
@@
This also highlights another alternative way to construct the probability mass function. Here, each component of the treated group’s shares before treatment is rescaled by the corresponding ratio of control group shares, and the resulting vector is then normalized to form a valid probability distribution.
@@ 2. Extensions With Multiple Time Periods
In this section, we propose several extensions of the method in the case where we have multiple time periods and possibly more potentially heterogeneous control groups.

### v2

@@ 1. Canonical $2\times 2$ DiD Case
In this section, I focus on the canonical, two-group, two-time period case in the DiD framework, and I clearly expose the assumptions and their implications. I will start with the analytical framework, later discuss the main identifying assumption, and all the economic and geometric implications.
1.1. Parallel growths and Identification
1.1.1. Analytical framework
In this section, I introduce the basic framework of my method. A categorical outcome $Y$ takes values in a finite set of $p$ mutually exclusive categories:
@@ 1.1.1. Analytical framework
I focus primarily on the absolute quantities for each category, which can appear in several empirical contexts. They can arise when the outcome is observed at the individual level. For example, in evaluating a job training program, individuals may be classified as unemployed, part-time employed, or full-time employed. The categorical quantity is obtained by aggregating individuals into these outcome categories. Sometimes, the data consist directly of quantities across categories rather than individual-level observations. For example, the number of votes each political party receives in an election. Unlike individual-level data, the identities or attributes of the units are not observed. In all these cases, the proposed method remains applicable, as it relies solely on knowledge of the categorical quantities.
I consider a setting with two groups, $g\in\{0,1\}$, where $g=1$ is the treated group and $g=0$ is the control group. These groups are observed in two time periods, $t\in\{0,1\}$, with $t=0$ being the pre-treatment period and $t=1$ the post-treatment period. The object of interest is a categorical outcome. For any category $c_{k}$, I define $q^{N}_{g,t}(c_{k})$: the potential untreated quantity (the quantity in $c_{k}$ if the group $g$ were untreated at time $t$). I also have $q^{I}_{g,t}(c_{k})$: the potential treated quantity (the quantity in $c_{k}$ if the group $g$ were treated at time $t$). For the untreated potential outcomes, I aggregate across categories to define the total quantity for a group-period:
$S^{N}_{gt}=\sum_{k=1}^{p}q^{N}_{g,t}(c_{k}).$
@@
All the same definitions apply to the treated potential quantities, $q^{I}_{g,t}(c_{k})$, and their associated totals and shares. From the data, I can identify the vectors $q^{N}_{0,0}$ for the control group and $q^{N}_{1,0}$ for the treated group in the pre-treatment period. In the post-treatment period, I can identify from data the untreated vector of quantities for the control group, $q^{N}_{0,1}$, and the treated counterpart for the treated group, $q^{I}_{1,1}$. The counterfactual of interest is the vector of quantities that I would have observed for the treated group in the absence of treatment $q^{N}_{1,1}$. Identifying this missing object is the core challenge of my analysis. Once this is identified, I can consider the treatment effect parameters.
I introduce two treatment effect parameters to capture the impact of the intervention. The first is the growth treatment effect on the treated (GTT), which quantifies the causal effect on the absolute size of a category. For each category $c_{k}$ (where $k=1,\ldots,q$), the GTT is defined as the proportional change in its quantity:
@@
The second parameter is the compositional treatment effect on the treated (CTT), which captures how treatment redistributes weight across categories. To measure this, I apply the compositional difference operator commonly used in compositional data analysis to compare probability mass functions. This operator quantifies how each category expands or contracts relative to all others. For intuition, fix a category $c_{k}$ and consider its proportional change:
@@
Dividing by the total change gives a vector showing each category’s relative gain or loss. I obtain the compositional difference operation used in compositional data analysis for the control group.
@@
1.1.2. Parallel growths
This section develops the identification strategy and clarifies its implications for the probability mass function. I start with the following assumption.
@@ 1.1.2. Parallel growths
This assumption ensures that all observed and counterfactual outcomes in the absence of treatment are defined over a common set of categories. It rules out situations where the support changes across groups or over time, which would make comparisons and counterfactual reasoning impossible. In practice, it implies that the set of possible outcomes is fixed. Such support conditions are standard in the literature. To introduce the main identifying assumption, I provide intuition for the DiD identification logic. The core idea is that, in the absence of treatment, the treated group would have followed the same trajectory as the control group. Figure 2 illustrates this DiD identification logic: the control group, unexposed to the treatment, serves as a counterfactual benchmark, mirroring what would have happened to the treated group had it not received the intervention.
@@
As I said before, in many empirical settings, the sizes of the groups defined in each category can vary dramatically in scale. Therefore, a standard parallel trends assumption on the raw quantities is problematic in this context, as it would impose the same absolute change on both large and small groups, which is often unrealistic. To address this issue of scale and to formulate a more plausible identifying assumption, I instead consider parallel trends on the log transformation of the quantities. This approach focuses on proportional, rather than absolute, changes. For a given vector, define the functions $(\log\cdot)$ and $(\exp\cdot)$ which apply respectively the logarithm and the exponential functions component-wise to the vectors.
@@
Assumption 2 is the core identifying condition for the model. It states that, in the absence of treatment, the proportional growth (or decline) of each category in the treated group would have been the same as that of the control group. This reflects the intuitive idea that underlying trends affect all categories proportionally to their size. This may appear as a more realistic and flexible assumption than one based on raw quantities in many economic contexts, especially when the sizes of categories differ significantly. It captures the idea that trends often affect all units proportionally to their size. It is also possible to use discrete covariates to make the assumption more believable. I discuss this possibility in appendix B. From this assumption, the counterfactual quantity vector for the treated group can be easily recovered, and the corresponding counterfactual probability mass can be obtained by normalization, as I see in Theorem 1. Taken together, assumptions 1 and 2 impose sufficient structure to point-identify the counterfactual distributions. The following theorem 1 formalizes this result
@@
This identification result is constructive and provides a straightforward way to implement the method. In the next section, I will discuss further justifications for parallel growths, exploring its economic and geometric implications.
1.2. Economic justifications of parallel growths
In this section, I provide economic justifications for parallel growths using the standard random utility framework (McFadden, 1972, 1974, 1977).
Consider a standard Random Utility Model, in which a decision-maker selects one alternative from a finite set of mutually exclusive choices $\{c_{1},c_{2},\dots,c_{p}\}$. The utility that the individual derives from alternative $c_{k}$ in the absence of treatment is modeled as:
$U^{N}_{gt}(c_{k})=V^{N}_{gt}(c_{k})+\varepsilon_{gt}(c_{k}),$
where $U^{N}_{gt}(c_{k})$ denotes the latent utility of category $k$ in group $g$ at time $t$ in the absence of treatment, and $V^{N}_{gt}(c_{k})$ is the systematic (deterministic) component of utility (can be seen as the measure of the total attractiveness of the category) and $\varepsilon_{gt}(c_{k})$ are the random unobserved shocks. The individual is assumed to choose the alternative that yields the highest latent utility:
$\text{Choice}=\arg\max_{c_{k}}U^{N}_{gt}(c_{k}).$
I decompose the total attractiveness into two parts:
$V^{N}_{gt}(c_{k})=\underbrace{\mu^{N}_{gt}(c_{k})-\log\left(\sum_{k}e^{\mu^{N}_{gt}(c_{k})}\right)}_{\text{relative attractiveness}}+\underbrace{\log\left(S^{N}_{gt}\right)}_{\text{absolute attractiveness}}.$
This decomposition shows that $V^{N}_{gt}(c_{k})$ contains two additive components: the first is the relative desirability of the category (within-choice structure) $\mu^{N}_{gt}(c_{k})-\log\Big(\sum_{j}e^{\mu_{gt}(c_{j})}\Big)$. This term reflects how desirable option $c_{k}$ is relative to all other alternatives at the same $(g,t)$. The second is absolute desirability (aggregate intensity) $\log(S_{gt})$. This term captures the overall size or participation level (e.g., how many people are making a choice at all). It scales all utilities equally, affecting the total mass of choices but not their relative shares. When it rises (say, due to higher purchasing power, stronger participation, or favorable aggregate conditions), all utilities shift up equally, people are more likely to choose something, but the proportions among categories might remain the same unless the first component also change. Together, they form a utility concept that blends relative preferences and absolute participation. This decomposition is useful to see how both components can be affected. We finally have:
$U^{N}_{gt}(c_{k})=\mu^{N}_{gt}(c_{k})-\log\left(\sum_{k}e^{\mu^{N}_{gt}(c_{k})}\right)+\log\left(S^{N}_{gt}\right)+\varepsilon_{gt}(c_{k}).$
In the traditional RUM, we assume that the random terms $(\varepsilon_{gt}(c_{1}),\dots,\varepsilon_{gt}(c_{p}))$ all follow standard type-1 extreme value distribution and are independent. Because each marginal error term remains Gumbel-distributed with mean $\gamma\approx 0.5772$ (the Euler–Mascheroni constant), the expected utility of alternative $c_{k}$ is:
$\mathbb{E}[U^{N}_{gt}(c_{k})]=\mu^{N}_{gt}(c_{k})-\log\left(\sum_{k}e^{\mu^{N}_{gt}(c_{k})}\right)+\log\left(S^{N}_{gt}\right)+\gamma.$
$\mathbb{E}[U^{N}_{gt}(c_{k})]$ captures the average latent total desirability in the population for category $c_{k}$. The first implication of the RUM is given by the following proposition.
Proposition 1 (Implication for expected utilities).
Under assumptions 1 and the RUM model assumptions, assumption  2 is equivalent to parallel trends of expected utilities:
$\mathbb{E}[U^{N}_{11}(c_{k})]-\mathbb{E}[U^{N}_{10}(c_{k})]=\mathbb{E}[U^{N}_{01}(c_{k})]-\mathbb{E}[U^{N}_{00}(c_{k})]\quad\forall\,c_{k}\in\bar{\mathcal{Y}}.$ (6)
This equivalence provides a clear economic interpretation of the identifying assumption. It states that, in the absence of treatment, both groups would have experienced the same evolution in how desirable each option is on average. For example, in a voting context, suppose the control group experiences a rise in the expected utility of voting Democrat (e.g., due to national political trends). My assumption implies that the treated group would have experienced the same increase in expected utility had it not been exposed to the policy. The same logic applies to all categories. Thus, the parallel growths assumption is not merely a mathematical condition on outcomes; it can also be seen as an economic hypothesis about the common evolution of underlying category-specific attractiveness in the absence of treatment. This provides a micro-founded justification for CoDiD: it estimates how a policy shifts the total attractiveness of alternatives as measured by expected utility.
Now, we will see how it also restrict the evolution of the shares. Define the multinomial logistic transform. For $\pi=[\pi(c_{1}),\dots,\pi(c_{p})]\in\mathcal{S}^{p-1}$. The log-odds (or multinomial logit) transformation $\ell:\mathcal{S}^{p-1}\to\mathbb{R}^{p-1}$ maps probabilities to real-valued indices:
@@ 1.2. Economic justifications of parallel growths
The parallel growths assumption implies a parallel trends assumption on the log-odds transformation of the probability distribution, as stated in this proposition.
@@
Assumptions 1 and 2 imply parallel trends in log-odds, i.e.,
$\ell(\pi_{1,1})-\ell(\pi_{1,0})=\ell(\pi_{0,1})-\ell(\pi_{0,0}).$
Moreover, parallel trends in log-odds is equivalent to parallel trends in utility differences for any pair of categories under the RUM:
$\Big\{\mathbb{E}[U^{N}_{11}(c_{k})]-\mathbb{E}[U^{N}_{11}(c_{l})]\Big\}-\Big\{\mathbb{E}[U^{N}_{10}(c_{k})]-\mathbb{E}[U^{N}_{10}(c_{l})]\Big\}=$
$\Big\{\mathbb{E}[U^{N}_{01}(c_{k})]-\mathbb{E}[U^{N}_{01}(c_{l})]\Big\}-\Big\{\mathbb{E}[U^{N}_{00}(c_{k})]-\mathbb{E}[U^{N}_{00}(c_{l})]\Big\},\quad\forall\,c_{k},c_{l}\in\bar{\mathcal{Y}}.$ (9)
The log-odds of choosing one category over another provide a direct empirical analogue to differences in the underlying expected utilities. They summarize how much more desirable one option is relative to another, capturing the relative strength or simply the preferences in the population. For instance, in a voting context, the log-odds of choosing Democrat over Republican quantify the population’s relative preference for Democrats. When the log-odds rise, it means that the expected utility, or average attractiveness, of voting Democrat has increased relative to that of voting Republican. The parallel growths assumption implies that, in the absence of treatment, these relative preferences would have evolved in parallel across groups for any pair of categories. That is, any secular change, such as a national political shift increasing the attractiveness of Democrats relative to Republicans, would have affected both treated and control groups identically had the treatment not occurred. The treatment effect is then identified as the deviation from this common trajectory. From that, I can also identify the counterfactual distribution with an alternative formula:
@@
In a $\times2$ DiD with categorical outcomes, I have shown that assuming parallel trends in log-quantity ratios is not ad hoc, it is equivalent to assuming parallel trends in expected utilities, a natural behavioral assumption under the Random Utility Model. Because utility differences (not levels) drive choice, this assumption is economically meaningful. This foundation justifies the CoDiD approach as a method for estimating how a policy changes the relative attractiveness of each alternative compared to others. In Appendix D, I extend this result to more general Random Utility Models that allow for dependence among the error terms, thereby relaxing the IIA assumption implicitly assumed in the standard multinomial RUM.
1.3. Geometric justifications of parallel growths
In this section, I provide some geometric justifications for parallel growths. I start by interpreting the restriction on the shares using the compositional difference operator.
Interpretation using the compositional difference interpretation: To further support the relevance of the assumption in this context, I provide an alternative interpretation that does not depend on the random utility framework or its underlying assumptions. Suppose our goal is to understand how the treatment redistributes probability mass across categories, that is, how each category gains or loses relative shares compared to the others as a consequence of treatment. An intuitive way of building the counterfactual would be to assume that the relative growth/decline in shares of each category in the control group reflects what would have happened in the treated group without treatment. The compositional difference operator $(\ominus)$ defined in equation 1 provides the precise mathematical language to formalize this intuition. I now show that the parallel growths assumption can be interpreted using that intuition. Specifically, assumption 2 implies that, in the absence of treatment, the relative growth or decline of category shares would have been similar across the treated and control groups. This interpretation is formally established in the following proposition.
@@ 1.3. Geometric justifications of parallel growths
In figure 3, each point in the triangle represents a distribution over three categories. As time evolves, distributions trace curves within the simplex. Although these paths may look non-linear in Euclidean space, in the simplex geometry, they are parallel, maintaining a consistent direction without intersecting. Here, I say that the implied evolution of the distributions is consistent with the nature of their space. Because of that, the compositional treatment effect on the treated can be written as:
@@
This expression mirrors the structure of the standard average treatment effect on the treated (ATT) in difference-in-differences, but it operates in the probability simplex rather than in the real line. In the classical DiD, differences in means capture changes in average outcomes across time and groups. In contrast, the compositional DiD captures differences in entire categorical distributions: each component of the treated group’s pre-treatment shares is rescaled by the corresponding ratio of control group shares, and the resulting vector is normalized to yield a valid probability distribution. Thus, (CTT) extends the ATT from the scalar outcome space to the compositional space, preserving the relative structure of category shares while identifying how treatment shifts the allocation of probability mass across categories. This also helps formalize the geometric implication of the assumption.
@@ 2. Extensions With Multiple Time Periods
In this section, I propose several extensions of the method in the case where we have multiple time periods and possibly more potentially heterogeneous control groups.